### v1

# Stabilizer codes for Heisenberg-limited many-body Hamiltonian estimation

Abstract
Estimating many-body Hamiltonians has wide applications in quantum technology. By allowing coherent evolution of quantum systems and entanglement across multiple probes, the precision of estimating a fully connected $k$-body interaction can scale up to $(n^{k}t)^{-1}$, where $n$ is the number of probes and $t$ is the probing time. However, the optimal scaling may no longer be achievable under quantum noise, and it is important to apply quantum error correction in order to recover this limit. In this work, we study the performance of stabilizer quantum error correcting codes in estimating many-body Hamiltonians under noise. When estimating a fully connected $ZZZ$ interaction under single-qubit noise, we showcase three families of stabilizer codes—thin surface codes, quantum Reed–Muller codes and Shor codes—that achieve the scalings of $(nt)^{-1}$, $(n^{2}t)^{-1}$ and $(n^{3}t)^{-1}$, respectively, all of which are optimal with $t$. We further discuss the relation between stabilizer structure and the scaling with $n$, and identify several no-go theorems. For instance, we find codes with constant-weight stabilizer generators can at most achieve the $n^{-1}$ scaling, while the optimal $n^{-3}$ scaling is achievable if and only if the code bears a repetition code substructure, like in Shor code.
I Introduction
High-precision measurement of parameters, or metrology, is crucial in all branches of science, especially in physics [1, 2, 3, 4, 5]. Classically, a common method to estimate an unknown parameter is to repeat the same experiment multiple times. As a consequence of the central limit theorem, the estimation error will scale as $1/\sqrt{t}$, known as the standard quantum limit (SQL), where $t$ is the time of experiment. Quantum mechanics allows us to achieve better precision by adopting the advantage of long-time coherent evolution, which results in an estimation error of order $1/t$, known as the Heisenberg limit (HL) [6, 7] which quadratically outperforms the SQL.
The HL, however, can be fragile against quantum noise [8, 9, 10, 11, 12]. To address this issue, quantum error correction (QEC) has been proposed as a tool to recover the HL under noise [13, 14, 15, 16, 11, 12, 17, 18, 19, 20]. Specifically, to estimate a Hamiltonian parameter under Markovian noise, it was known the HL is achievable if and only if the “Hamiltonian-not-in-Lindblad-span” (HNLS) condition is satisfied [11, 12], in which case a QEC protocol was proposed to recover the HL. Despite the generality of the QEC protocol, its implementation can be challenging in practice, partially due to the lack of efficient encoding and decoding procedure, and the requirement of a noiseless ancillary system as part of the code [12, 19]. Using stabilizer codes [21, 22], a class of QEC codes with natural encoding and decoding methods via stabilizer measurements and no requirement of noiseless ancilla, for quantum metrology can alleviate the above constraints. So far, stabilizer codes have rarely been explored for quantum metrology in previous works [12, 17, 18, 19, 20], except for repetition codes for phase estimation against bit-flip noise [13, 14, 15].
In multi-probe systems, estimating the strength of Hamiltonians has always been a major topic in quantum metrology [23, 24, 25, 26, 6, 7, 27, 28, 29, 30]. The scaling of the estimation error with respect to the number of probes $n$ can achieve up to $1/n$ for 1-local Hamiltonians, using entangled states, e.g., the Greenberger–Horne–Zeilinger (GHZ) state [7] and we will refer to it as the Heisenberg scaling (HS) [^1] Note that the scaling $1/n$ can also sometimes be referred to as the Heisenberg limit (with respect to the number of probes). However, in this paper we call it the HS to distinguish it from the HL with respect to probing time $t$..
The estimation with higher-order interactions can go beyond the HS [30, 31, 32]. For example, estimating a $k$-local Hamiltonian can reach a precision of order $1/n^{k}$, usually referred to as the super-Heisenberg scaling (SHS) when $k>1$ [30]. Similar to the HL, the HS and SHS are also compromised by the effect of quantum noise. For example, the HS for estimating 1-local Hamiltonians will degrade to $1/\sqrt{n}$ under generic single-qubit noise [10, 9, 11, 12]. In general, it is of great interest to explore which types of SHS are recoverable for estimating many-body Hamiltonians under different types of quantum noise.
In this work, we consider the achievability of the HL, HS and SHS for estimating a fully connected $ZZZ$ interaction under single-qubit noise using stabilizer codes. First, we will present three classes of $[[n,1,3]]$ stabilizer codes—thin surface codes [33, 34, 35], quantum Reed-Muller codes [36] and Shor codes [37]—that achieve the HL and HS. This demonstrates the applicability of stabilizer codes in noisy quantum metrology. In particular, the optimal SHS $1/n^{3}$ can be achieved by Shor codes. Secondly, we will dive into the achievability of the SHS and explore the roles of different code structures, e.g., non-locality of stabilizer generators, repetition code substructures, etc., that are necessary for achieving different types of SHS. One prominent result is the non-achievability of the SHS when stabilizer generators have constant weights, e.g., for quantum low-density parity-check (LDPC) codes [38]. These no-go results then imply the optimality of the previously mentioned three classes of codes under different constraints on code structures, and may further provide guidance on identifying or constructing stabilizer codes that are suitable for many-body Hamiltonian estimation.
II Formulation of the problem
Hamiltonian evolution under Markovian noise can be described by the Lindblad master equation [39, 40, 41]
$\frac{d\rho}{dt}=-i[H,\rho]+\sum_{a}\left(L_{a}\rho L_{a}^{\dagger}-\frac{1}{2}\{L_{a}^{\dagger}L_{a},\rho\}\right),$ (1)
where $\rho$ is the density matrix of the probe system, $H$ is the Hamiltonian and $\{L_{a}\}_{a}$ are the Lindblad operators describing quantum noise. We consider the situation where the probe system consists of $n$ qubits ($n\geq 3$),
$H=\omega\sum_{1\leq i<j<k\leq n}Z_{i}Z_{j}Z_{k}+H_{0}\coloneqq\omega G+H_{0},$ (2)
where $H_{0}$ is any (unknown) 2-local Hamiltonian, $\omega$ is the strength of the 3-local Hamiltonian $G$ that we want to estimate, and $Z_{i}$ represents the Pauli-Z operator acting on the $i$-th qubit (we will use $X,Y,Z$ to denote Pauli-X,Y,Z operators, respectively). The set $\{L_{a}\}_{a}$ consists of local operators (i.e., each $L_{a}$ acts non-trivially on only one qubit). We further assume that the noise is complicated enough that the Lindblad span
${\mathcal{S}}$ :-$\coloneq{\rm span}_{\rm H}\{I,L_{a},L_{a^{\prime}}^{\dagger},L_{a}^{\dagger}L_{a^{\prime}}\}_{\forall{a,a^{\prime}}}:$ (3)
$={\rm span}_{\rm H}\{P_{i}P_{j}\}_{\forall{i,j;P\in\{X,Y,Z,I\}}}$ (4)
[…]
Consider the quantum metrological protocol where an initial $n$-qubit state $\rho(0)$ is prepared, quantum controls are applied sufficiently fast during the evolution, and $\omega$ is estimated through quantum measurement on the output state $\rho_{\omega}(t)$ at time $t$ [16, 11, 12]. The estimation precision $\delta\omega$ can be given by the quantum Cramér–Rao bound [42, 43]
$\delta\omega\geq 1/\sqrt{\nu F(\rho_{\omega}(t))},$ (5)
where $\nu$ is the number of repeated experiments and $F(\rho_{\omega}(t))$ is the quantum Fisher information (QFI) of the state $\rho_{\omega}(t)$ [44, 45]. The quantum Cramér–Rao bound is saturable as $\nu\rightarrow\infty$ [46], and thus the scaling of $F(\rho_{\omega}(t))$ determines the scaling of $\delta\omega$ with respect to $t$ and $n$. Specifically, we say that the HL is achieved when $\delta\omega\sim t^{-1}$ or $F(\rho_{\omega}(t))=\Theta(t^{2})$, the HS is achieved when $\delta\omega\sim n^{-1}$ or $F(\rho_{\omega}(t))=\Theta(n^{2})$, and the SHS is achieved when $\delta\omega\sim n^{-k}$ or $F(\rho_{\omega}(t))=\Theta(n^{2k})$ for any $k>1$.
In the noiseless case ($L_{a}=H_{0}=0$), the optimal initial state is the GHZ state
|$\ket{\psi(0)}=\frac{1}{\sqrt{2}}({\ket{0^{\otimes n}}+\ket{1^{\otimes n}}}),$ (6)
which leads to an output state
|$\ket{\psi_{\omega}(t)}=\frac{e^{i\omega\frac{n(n-1)(n-2)}{6}}\ket{0^{\otimes n}}+e^{-i\omega\frac{n(n-1)(n-2)}{6}}\ket{1^{\otimes n}}}{\sqrt{2}},$ (7)
after evolution time $t$, and the corresponding QFI of the output state $\rho_{\omega}(t)=\ket{\psi_{\omega}(t)}\bra{\psi_{\omega}(t)}$ | is
$F(\rho_{\omega}(t))=4\left(\frac{n(n-1)(n-2)}{6}\right)^{2}t^{2}=\Theta(n^{6}t^{2}),$ (8)
which achieves the HL and the SHS.
In general, the HNLS condition [12] states that it is possible to achieve the HL if and only if $G\not\in{\mathcal{S}}$, which is satisfied in our setting. The QEC protocol utilizes a QEC code that perfectly corrects all single-qubit errors, i.e., $\Pi S\Pi\propto\Pi$ for all $S\in{\mathcal{S}}$, where $\Pi$ is the projection onto the code subspace [47, 48]. In the fast control limit (i.e., fast quantum controls can be applied at an arbitrarily high frequency), the evolution of $\rho_{\textsc{l}}=\Pi\rho\Pi$ is given by
$\frac{d\rho_{\textsc{l}}}{dt}=-i\omega[G_{\mathrm{eff}},\rho_{\textsc{l}}],$ (9)
where $G_{\mathrm{eff}}=\Pi G\Pi\not\propto\Pi$. (Note that in the noiseless case, $\Pi=I$ and the discussion below also holds.) The QFI is maximized when the initial state is chosen as |$\ket{\psi(0)}=\frac{1}{\sqrt{2}}{\ket{\lambda_{\min}}+\ket{\lambda_{\max}}}$, that leads to
$F(\rho_{\omega}(t))=(\lambda_{\max}-\lambda_{\min})^{2}t^{2}\eqcolon(\Delta G_{\mathrm{eff}})^{2}t^{2},$ (10)
which achieves the HL, where $\lambda_{{\min},{\max}}$ are the smallest (and largest) eigenvalues that correspond to the eigenstates |$\ket{\lambda_{{\min},{\max}}}$ of $G_{\mathrm{eff}}$ that are restricted in the code subspace. $\Delta G_{\mathrm{eff}}:=\lambda_{\max}-\lambda_{\min}$. Note that the above implies that it is sufficient to consider a two-level logical system, as we can always, without loss of generality, restrict the code subspace to ${\rm span}\{\ket{\lambda_{{\min}}},\ket{\lambda_{{\max}}}\}$.
Among all possible ancilla-assisted QEC codes, the optimal QEC protocol [12] can achieve $(\Delta G_{\mathrm{eff}})_{{\mathrm{opt}}}=2\left\|G-{\mathcal{S}}\right\|$, where $\left\|G-{\mathcal{S}}\right\|$ is the distance between $G$ and ${\mathcal{S}}$ in terms of the operator norm. Specifically in our case (see Appx. A),
$(\Delta G_{\mathrm{eff}})_{{\mathrm{opt}}}=2\!\min_{\text{2-local }\!S}\left\|G-S\right\|=\frac{n^{3}}{12}+O(n^{2}).$ (11)
Comparing this to the noiseless case, where
$\Delta G=\frac{n^{3}}{3}+O(n^{2}),$ (12)
we see that the optimal QFI for estimating $ZZZ$ interaction under generic qubit noise is $F(\rho_{\theta}(t))=\Theta(n^{6}t^{2})$, which bears the same scaling as the noiseless one.
Although QEC can in principle recover the HL and the SHS for estimating the 3-local Hamiltonian $G$ from the above discussion, it is unclear how the QEC protocol (with encoding, decoding and measurement operations) can be implemented in reality. In this work, we consider stabilizer QEC for the task of estimating $ZZZ$ interactions, which is the most popular family of QEC and was demonstrated recently in experiments across multiple platforms [49, 50, 51]. One advantage of it is the standard encoding, decoding and measurement procedures via syndrome measurements [21] that were currently unavailable in the known QEC protocols for metrology. In addition, as stabilizer codes have already been theoretically extensively studied in previous literature for quantum memory, quantum computation, etc., it is natural to explore their applications in other noisy information processing tasks like quantum metrology and identify the code properties that are crucial for many-body Hamiltonian estimation.
III Achieving the HL, HS and SHS
As discussed in the previous section, the HL is achievable using a QEC code if and only if $G_{\mathrm{eff}}=\Pi G\Pi$ acts non-trivially in the code subspace. In this section, we first establish a relation between $\Delta G_{\mathrm{eff}}$ and the logical operators of the stabilizer code, and then present three families of stabilizer codes that achieve both the HL and the HS (or SHS).
III.1 Relation between QFI and $Z^{\otimes 3}$-type Logical Operators
Here we use the stabilizer formalism to establish a relation between the QFI and the number of $Z^{\otimes 3}$-type logical operators, i.e., operator $Z_{i}Z_{j}Z_{k}$ for some $i,j,k$ that maps logical states to logical states in a non-trivial way, and this relation will be used throughout the remainder of the paper.
Lemma 1.
Consider using an $[[n,1,3]]$ [^2] We use $[[n,k]]$ ($[[n,k,d]]$) to represent a stabilizer code that encodes $k$ logical qubits into $n$ physical qubits (with distance $d$). stabilizer code to estimate $\omega$ in Eq. (1) and Eq. (2). Let $\ell$ be the number of $Z_{i}Z_{j}Z_{k}$ operators appearing in $G$ that are logical operators of the code. Then the output QFI is
$F(\rho_{\omega}(t))=4\ell^{2}t^{2}.$ (13)
Proof.
Let ${\mathscr{S}}$ and ${\mathscr{N}}({\mathscr{S}})$ be the stabilizer and normalizer sets of the stabilizer code, respectively. A term $Z_{i}Z_{j}Z_{k}$ of $G$ belongs to one of the following three classes: (1) $Z_{i}Z_{j}Z_{k}\in{\mathscr{S}}$, (2) $Z_{i}Z_{j}Z_{k}\not\in{\mathscr{N}}({\mathscr{S}})$, and (3) $Z_{i}Z_{j}Z_{k}\in$ ${\mathscr{N}}({\mathscr{S}})\setminus{\mathscr{S}}$. The operator is called a logical operator if it belongs to class (3). By Eq. (10), the QFI is proportional to $\Delta G_{\mathrm{eff}}$ where $G_{\mathrm{eff}}=\Pi G\Pi=\sum_{ijk}\Pi Z_{i}Z_{j}Z_{k}\Pi$ with $\Pi$ being the projection onto the code space. We separately calculate the contributions of the three types of terms to the $\Delta G_{\mathrm{eff}}$.
  1. (1)
    Suppose $Z_{i}Z_{j}Z_{k}\in{\mathscr{S}}$. It immediately follows that $Z_{i}Z_{j}Z_{k}\Pi=\Pi$ such that $\Pi Z_{i}Z_{j}Z_{k}\Pi=\Pi$. Hence, the terms that are in ${\mathscr{S}}$ do not contribute to $\Delta G_{\mathrm{eff}}$.
  2. (2)
    Suppose that $Z_{i}Z_{j}Z_{k}\not\in{\mathscr{N}}({\mathscr{S}})$. Then, there is an operator $Q\in{\mathscr{S}}$ that anti-commutes with $Z_{i}Z_{j}Z_{k}$. So, $\Pi Z_{i}Z_{j}Z_{k}\Pi=\Pi QZ_{i}Z_{j}Z_{k}\Pi=-\Pi Z_{i}Z_{j}Z_{k}Q\Pi=-\Pi Z_{i}Z_{j}Z_{k}\Pi$, which leads to $\Pi Z_{i}Z_{j}Z_{j}\Pi=0$. Hence, the terms that are not in the normalizer group also do not contribute to the QFI.
  3. (3)
    Suppose that $Z_{i}Z_{j}Z_{k}\in{\mathscr{N}}({\mathscr{S}})\setminus{\mathscr{S}}$. As all the $ZZZ$ terms commute with each other, the non-trivial logical operators must be different representatives of the same logical operator. In addition, $\Pi Z_{i}Z_{j}Z_{k}\Pi$ must correspond to a logical Pauli operator in the two-dimensional code subspace, which means $\Delta(\Pi Z_{i}Z_{j}Z_{k}\Pi)=2$.
Hence, if the number of $Z^{\otimes 3}$-type logical operators in $G$ is $\ell$, the QFI is given by $F(\rho_{\omega}(t))=(\Delta G_{\mathrm{eff}})^{2}t^{2}=4\ell^{2}t^{2}$. ∎
Lemma 13 implies that the HL is achievable for any stabilizer codes with $Z^{\otimes 3}$-type logical operators. The scaling of the QFI with respect to $n$ is determined by the scaling of $\ell$ with respect to $n$. In the following parts of this section, we will apply Lemma 13 to demonstrate the achievability of the HS ($F(\rho_{\omega}(t))=\Theta(n^{2})$) and SHS ($F(\rho_{\omega}(t))=\Theta(n^{4}),\Theta(n^{6})$) using three different families of stabilizer codes, or more specifically, Calderbank–Shor–Steane (CSS) codes [52, 53].
III.2 Thin Surface Code
First, we consider a special kind of surface code [33, 34, 35] that we will refer to as the thin surface code, which corrects single-qubit $Z$ errors. Fig. 1 shows the $[[n,1,3]]$ thin surface code with $n=33$. The qubits are placed on the edges of a $(2\times(n-3)/5)$ lattice. On the primary (solid) lattice, the $X$-type stabilizer generators are $X^{\otimes 4}$ plaquette operators in the bulk and on the smooth edges, and $X^{\otimes 3}$ plaquette operators on the rough edges. Similarly, the $Z$-type stabilizer generators are $Z^{\otimes 4}$ and $Z^{\otimes 3}$ plaquette operators on the dual (dashed) lattice,
Figure 1: $[[33,1,3]]$ thin surface code. Physical qubits are placed on the edges of the solid lattice (including the start and end points of the horizontal solid lines). $X$/$Z$-type stabilizer and logical operators are represented by solid/dashed lines that pass through the involved qubits.
The quantum code encodes one logical qubit into $n$ physical qubits. A logical $X$ operator is a string of $X$ operators connecting the two rough edges on the primary lattice. Similarly, a logical $Z$ operator is a string of $Z$ operators connecting the two smooth edges on the dual lattice.
For the $[[n,1,3]]$ thin surface code, the $Z^{\otimes 3}$-type logical operators must correspond to one vertical $Z^{\otimes 3}$ strings on the dual lattice. This is because multiplying any one of them by a stabilizer can either increase its weight, or transform it to another $Z^{\otimes 3}$-type logical operator. By a simple counting, the number of such $Z^{\otimes 3}$ strings for an $n$-qubit code is $(n+2)/5$. Therefore, by Lemma 13, we have the following result.
Theorem 1.
The $[[n,1,3]]$ thin surface code achieves a QFI of ${4(n+2)^{2}t^{2}}/{25}=\Theta(n^{2}t^{2})$.
III.3 Quantum Reed-Muller Code
The stabilizer group of a CSS code is generated by operators from two subsets ${\mathsf{S}}_{x}$ and ${\mathsf{S}}_{z}$, where the subset ${\mathsf{S}}_{x}$ (${\mathsf{S}}_{z}$) consists of stabilizer generators that are tensor products of $X$ ($Z$) and $I$. Here, we consider quantum Reed–Muller (QRM) codes that are CSS codes whose ${\mathsf{S}}_{x}$ and ${\mathsf{S}}_{z}$ are obtained from the parity-check matrices of two classical Reed–Muller (RM) codes [36, 54].
We first review the definition of classical RM codes [55]. Define $\bm{v}_{0}^{T}=(1,1,\ldots,1)\eqcolon\bm{1}_{2^{m}}^{T}$ as the $2^{m}$-tuple row vector consisting entirely of 1s. Throughout this section, we will use the notation $\bm{0}_{p}$ to denote a vector of $p$ zeros, and $\bm{1}_{p}$ to denote a vector of $p$ ones. For any integer $i$ with $1\leq i\leq m$, define the $2^{m}$-tuple row vector $\bm{v}_{i}^{T}=(\bm{0}_{2^{i-1}}^{T}|\bm{1}_{2^{i-1}}^{T}|\bm{0}_{2^{i-1}}^{T}|\ldots|\bm{1}_{2^{i-1}}^{T})$, where “||\,|\,|” denotes the concatenation of row vectors. The generator matrix of $RM(1,m)$, i.e., the 1st order RM code of length $2^{m}$, is given by
${\mathsf{G}}(1,m)=\begin{pmatrix}\text{---}\bm{v}_{0}^{T}\text{---}\\ \text{---}\bm{v}_{1}^{T}\text{---}\\ \vdots\\ \text{---}\bm{v}_{m}^{T}\text{---}\\ \end{pmatrix}.$ (14)
The $r$-th order RM code, denoted by $RM(r,m)$, is constructed by taking the Boolean products (denoted by *) of the codewords of a 1st-order RM code. In particular, the generator matrix ${\mathsf{G}}(r,m)$ of $RM(r,m)$ is obtained by taking the rows of $RM(1,m)$ along with all the distinct $2,3,\ldots,r$-fold Boolean products of the rows of ${\mathsf{G}}(1,m)$. For example, the set of rows of the generator matrix ${\mathsf{G}}(2,3)$ is $\{\bm{v}_{0}^{T},\bm{v}_{1}^{T},\bm{v}_{2}^{T},\bm{v}_{3}^{T},\bm{v}_{1}^{T}*\bm{v}_{2}^{T},\bm{v}_{1}^{T}*\bm{v}_{3}^{T},\bm{v}_{2}^{T}*\bm{v}_{3}^{T}\}$, where ${\mathsf{G}}(1,3)$ corresponds to the first four rows. It can be shown that $RM(r,m)$ is a $[2^{m},N(m,r),2^{m-r}]$ code, where $N(m,r)\coloneqq 1+\binom{m}{1}+\binom{m}{2}+\cdots+\binom{m}{r}$ and $RM(r,m)^{\perp}=RM(m-r-1,m)$ [55].
The QRM codes are derived from shortened RM codes $\overline{RM}(r,m)$ whose generating matrix $\overline{{\mathsf{G}}}(r,m)$ is obtained by deleting the first row and the first column of ${\mathsf{G}}(r,m)$. The following three properties of a shortened RM code $\overline{RM}(r,m)$ will be useful to us. Their proofs are given in Appx. B.
  1. (1)
    $\overline{RM}(1,m)$ is strictly contained in $\overline{RM}(m-2,m)$.
  2. (2)
    The vector $\bm{1}_{2^{m}-1}$ and those corresponding to the rows of $\overline{{\mathsf{G}}}(m-2,m)$ form a basis for $\overline{RM}(1,m)^{\perp}$.
  3. (3)
    The number of codewords in $\overline{RM}(1,m)^{\perp}$ with weight $(2^{m}-4)$ is $\frac{1}{6}(4^{m}-3\cdot 2^{m}+2)$.
Here, we focus on $QRM(1,m)$ which is an $[[n=2^{m}-1,1,3]]$ code (see a definition of general $QRM(r,m)$ in e.g., [56]). The set of $X$-type stabilizer generator ${\mathsf{S}}_{x}$ is obtained from the rows of $\overline{{\mathsf{G}}}(1,m)$ by substituting $1$s by $X$s and 00s by $I$s in the tensor product forms of the operators. The elements of ${\mathsf{S}}_{z}$ are similarly obtained from $\overline{{\mathsf{G}}}(m-2,m)$ by replacing $1$s by $Z$s and 00s by $I$s. Due to Property (1), the basis codewords of the CSS code $QRM(1,m)$ are given by |$\ket{z+\overline{RM}(1,m)}=\frac{1}{2^{m}}\sum_{y\in\overline{RM}(1,m)}\ket{z+y}$ for $z\in\overline{RM}(m-2,m)$ [22]. Note that $Z^{\otimes n}$ commutes with all the $X$-type stabilizer generators since each generator is of even weight; however, $Z^{\otimes n}$ is not in the stabilizer group by Property (2). Thus, $Z^{\otimes n}$ is a logical operator and one can choose a specific basis in the logical subspace such that it is a logical $Z$ operator, i.e.,
$\overline{Z}:=Z^{\otimes n}=Z^{\otimes(2^{m}-1)}.$ (15)
Following Lemma 13, we obtain the scaling of the QFI with respect to $n$ by calculating the number of weight-3 representatives of the logical Pauli-Z operator.
Theorem 2.
The $[[n=2^{m}-1,1,3]]$ $QRM(1,m)$ code achieves a QFI of ${n^{2}(n-1)^{2}}t^{2}/{9}=\Theta(n^{4}t^{2})$.
Proof.
Operators equivalent to the logical $Z$ operator in Eq. (15) are obtained by multiplying $\overline{Z}$ with different stabilizers. Since the weight of $\overline{Z}$ is $12^{m}-$ 1, it follows that the number of $Z^{\otimes 3}$-type logical operators $\ell$ is the same as the number of $Z^{\otimes(n-3)}=Z^{\otimes(2^{m}-4)}$ stabilizers. This amounts to calculating the number of weight-$(2^{m}-4)$ codewords in the shortened code $\overline{RM}(m-2,m)$.
Below we prove the number of weight-$(2^{m}-4)$ codewords in $\overline{RM}(m-2,m)$ is equal to that in $\overline{RM}(1,m)^{\perp}$. Using Property (3), this implies that $\ell=\frac{1}{6}(4^{m}-3\cdot 2^{m}+2)=\frac{1}{6}n(n-1)$. Combining with Lemma 13, the theorem is then proven.
To prove the statement above, let $V=\{\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{q}\}$ be the set of weight-$(2^{m}-4)$ vectors in $\overline{RM}(m-2,m)$. Using Property (2), it follows that $V\cup\{\bm{1}_{n}+\bm{u}_{i}:1\leq i\leq p\}$ is the set of weight-$(2^{m}-4)$ vectors in $\overline{RM}(1,m)^{\perp}$, where $U=\{\bm{u}_{1},\bm{u}_{2},\ldots,\bm{u}_{p}\}$ is the set of weight-3 vectors in $\overline{RM}(m-2,m)$. We know that the weights of the codewords of the unshortened $RM(m-2,m)$ are even [^3] Can be shown using the Macwilliams identity [55] (see Appx. B).. Therefore, the weight-3 codewords of $\overline{RM}(m-2,m)$ are exactly the shortened versions of the weight-4 codewords of $RM(m-2,m)$ that have 1 as the first entry. So, for each $\bm{u}_{i}\in U$, we have a distinct weight-4 vector $(1|\bm{u}_{i}^{T})^{T}\in RM(m-2,m)$. This vector is of the form $(1|\bm{u}_{i}^{T})^{T}=\bm{1}_{2^{m}}+\bm{w}$, where $\bm{w}\in RM(m-2,m)$ is a weight-$(2^{m}-4)$ vector whose first entry is 0. Hence, $\bm{w}^{T}=(0|\bm{v}_{j}^{T})$ for some $\bm{v}_{j}\in V$. This implies that $\bm{v}_{j}=\bm{1}_{n}+\bm{u}_{i}$. Therefore, $V\cup\{\bm{1}_{n}+\bm{u}_{i}:1\leq i\leq p\}=V$, thus completing the proof. ∎
The QRM code achieves the SHS, which outperforms the thin surface code. As we will see, we can achieve a better scaling using Shor codes.
III.4 Shor Code and Generalization
We now consider the $n$-qubit Shor code, which is a simple generalization of the $9$-qubit Shor code [37], where $n=3n_{r}$ is divisible by $3$. The code is constructed by concatenating a phase flip code (|$\left(\frac{\ket{0}\pm\ket{1}}{\sqrt{2}}\right)^{\otimes 3}$ of distance three with an $n_{r}$-qubit repetition code. (Note that here a concatenation of code 1 and code 2 means encoding every physical qubit of code 1 using code 2.) The logical basis states are given by
|$\ket{0_{\textsc{l}}}=\frac{1}{\sqrt{8}}\left(\ket{0}^{\otimes n_{r}}+\ket{1}^{\otimes n_{r}}\right)^{\otimes 3},$ (16a)
|$\ket{1_{\textsc{l}}}=\frac{1}{\sqrt{8}}\left(\ket{0}^{\otimes n_{r}}-\ket{1}^{\otimes n_{r}}\right)^{\otimes 3}.$ (16b)
A particularly nice set of stabilizer generators for the code is given in Table 1. It is clear that one logical qubit is encoded into $n=3n_{r}$ physical qubits.
    Name Operator
    $g_{1}$ $Z$ $Z$
    $g_{2}$ $Z$ $Z$
    $\vdots$ $\ddots$
    $g_{n_{r}-1}$ $Z$ $Z$
    $g_{n_{r}}$ $Z$ $Z$
    $\vdots$ $\ddots$
    $g_{2n_{r}-2}$ $Z$ $Z$
    $g_{2n_{r}-1}$ $Z$ $Z$
    $\vdots$ $\ddots$
    $g_{3n_{r}-3}$ $Z$ $Z$
    $g_{3n_{r}-2}$ $X$ $X$ $X$cdots $X$ $X$ $X$ $X$cdots $X$ $X$
    $g_{3n_{r}-1}$ $X$ $X$cdots $X$ $X$ $X$ $X$cdots $X$ $X$
Table 1: A set of stabilizer generators of the Shor code. The code is divided into three blocks of length $n_{r}$. Each block supports $(n_{r}-1)$ $Z$-type stabilizers, and two consecutive $Z$-type stabilizers overlap on one qubit within each block. Two $Z$-type stabilizers from different blocks don’t overlap. There are two $X$-type stabilizer generators $g_{3n_{r}-2}$ and $g_{3n_{r}-1}$, each with support on a string of $2n_{r}$ qubits. The supports of the two $X$-type generators overlap on the middle block of qubits.
One set of representatives of the logical $X$ and logical $Z$ operators is
$\overline{X}$ $:=\left(Z\otimes I^{\otimes(n_{r}-1)}\right)^{\otimes 3},$ (17a)
$\overline{Z}$ $:=X^{\otimes{n}}.$ (17b)
Note that the logical operator $\overline{X}$ in Eq. (17a) is a term in $G=\sum_{i<j<k}Z_{i}Z_{j}Z_{k}$. Thus, all $Z^{\otimes 3}$-type logical operators must be representatives of the logical $X$ operator. From the definition of the logical basis states in Eq. (16), it is clear that $Z_{i}Z_{j}Z_{k}$ is a logical operator if and only if the three $Z$ operators are supported on different blocks. There are $n_{r}^{3}=(n/3)^{3}$ such logical operators. Using Lemma 13, we have the following result.
Theorem 3.
The $[[n,1,3]]$ Shor code achieves a QFI of ${4n^{6}t^{2}}/{27}=\Theta(n^{6}t^{2})$.
Inspired by the repetition code substructure in the Shor code, we have a more general statement:
Theorem 4.
Let an $[[n,1]]$ code $C$ be the concatenation of a $[[q,1]]$ stabilizer code $C^{\prime}$ that has at least one $Z^{\otimes 3}$-type logical operator and can correct single-qubit Pauli-Z errors and the $[[n/q,1]]$ repetition code, where $q$ is a constant integer and $n$ is divisible by $q$. Then $C$ achieves a QFI of $\Theta(n^{6}t^{2})$.
Proof.
We need to prove (1) $C$ has $\Theta(n^{3})$ $Z^{\otimes 3}$-type logical operators and (2) $C$ is an $[[n,1,3]]$ stabilizer code.
Suppose that $Z_{1}Z_{2}Z_{3}$ is a logical operator for $C^{\prime}$, it then implies $Z_{i}Z_{j}Z_{k}$ are logical operators for $C$ whenever $1\leq i\leq n_{r}$, $n_{r}+1\leq j\leq 2n_{r}$, $2n_{r}+1\leq k\leq 3n_{r}$ where $n_{r}:=n/q$. Note that in $C$, qubits with labels from $q_{1}n_{r}+1$ to $(q_{1}+1)n_{r}$ correspond to the $q_{1}$-th qubit in $C^{\prime}$ expanded from the repetition code concatenation for any $q_{1}$. The number of different $Z_{i}Z_{j}Z_{k}$ above are $n_{r}^{3}=\Theta(n^{3})$, proving statement (1).
To prove statement (2), we need to show $C$ can correct any single-qubit error, i.e., any operator $S_{i}S_{j}$ for $S=I,X,Z$ must either belong to the stabilizer group of $C$ or anti-commute with one stabilizer of $C$. First, note that for any $i$, $X_{i}$ must anti-commute with $Z_{i}Z_{k}$ for any $k\neq i$ and $\lceil i/n_{r}\rceil(n_{r}-1)+1\leq k\leq\lceil i/n_{r}\rceil n_{r}$ which is a stabilizer of $C$. It implies any $S_{i}S_{j}$ anti-commutes with some stabilizer of $C$ whenever one of them is a Pauli-X operator. On the other hand, any $n$-qubit operator $S_{i}S_{j}$ with $S=I,Z$ can be mapped to a $q$-qubit operator $S_{\lceil i/n_{r}\rceil}S_{\lceil j/n_{r}\rceil}$ acting on $C^{\prime}$ which either belongs to the stabilizer group of $C^{\prime}$ or anti-commutes with one stabilizer of $C^{\prime}$ because $C^{\prime}$ corrects single-qubit Pauli-Z errors. It implies the required condition is also satisfied for $S_{i}S_{j}$ with $S=I,Z$. Statement (2) is then proven. ∎
Theorem 3 is a special case of Theorem 4 where $C^{\prime}$ is the three-qubit phase-flip code. As discussed in Sec. II, the above codes provide the optimal SHS scaling we can achieve for estimating the $ZZZ$ interaction under single-qubit noise (in the setting of Eq. (1) and Eq. (2)).
IV Necessary conditions for achieving the SHS
In the previous section, we discussed three examples of quantum codes that asymptotically achieve $(nt)^{-1}$, $(n^{2}t)^{-1}$, and $(n^{3}t)^{-1}$ precision scalings for the $ZZZ$ interaction estimation problem. The three families of codes have special stabilizer structures: the thin surface code has a set of low-weight stabilizer generators; the QRM code lacks a complete set of low-weight stabilizer generators; and the Shor code has local low-weight $Z$-type generators even though the $X$-type generators scaled linearly with the length of the code. In this section, we address the question of what properties a stabilizer code must have to achieve different precision scalings with respect to $n$, or equivalently, different scalings of the number of $Z^{\otimes 3}$-type logical operators with respect to $n$.
IV.1 Unachievability of the SHS
We proved in Theorem 1 that the thin surface code can be used to achieve $1/n$ precision scaling in $\omega$. In fact, this scaling can more generally apply to stabilizer codes with constant weight generators that can correct arbitrary single qubit errors. We want to find an upper bound on the number of $Z^{\otimes 3}$-type logical operators for such codes, which will consequently give us the best possible precision scaling.
Theorem 5.
Let $C$ be an $[[n,1,3]]$ stabilizer code such that there exists a set of stabilizer generators with weights at most $w$. Then, $\ell$, the number of $Z^{\otimes 3}$-type logical operators of $C$ is at most $32w(w+1)n/$ 3. In particular, if $w$ is a constant independent of $n$, then $\ell=O(n)$.
Proof.
Let ${\mathsf{S}}$ be a stabilizer generator set of $C$ such that the weight of every $g\in{\mathsf{S}}$ satisfies $\operatorname{wt}(g)\leq w$. Without loss of generality, we assume there are no $Z_{i}$-type stabilizers. Otherwise, we can always remove from the entire set of $n$ qubits the set of qubits on which a single Pauli-Z are stabilizers, i.e., qubits that are always in states |$\ket{0}$. Then the remaining $\tilde{n}$ qubits give an $[[\tilde{n},1,3]]$ stabilizer code $\tilde{C}$ satisfying the same weight constraint without $Z_{i}$-type stabilizers. Any $Z^{\otimes 3}$-type logical operator in $C$ must be fully supported on the remaining $\tilde{n}$ qubits and is thus also a $Z^{\otimes 3}$-type logical operator in $\tilde{C}$. Then by proving the theorem for $\tilde{C}$, the theorem for $C$ also holds.
We define the degree of a qubit as the number of $Z^{\otimes 3}$-type logical operators that are supported on it. Without loss of generality, assume that the first qubit (i.e., the qubit labeled by $1$) has the highest degree. Define a graph ${\mathfrak{G}}$ with $(n-1)$ vertices corresponding to the qubits labeled by $2,3,\ldots,n$. There is an edge between the $i$-th and $j$-th vertices if $Z_{1}Z_{i}Z_{j}$ is a logical operator of the code. We define the order of a vertex as the number of edges connected to it.
Consider the two separate cases: (1) There is some vertex with order $>w$, and (2) The order of every vertex is $\leq w$.
Case 1: Suppose that the order of vertex $i_{1}$ is greater than $w$. Then, there are vertices $i_{2},i_{3},\ldots,i_{w+2}$ that are connected to vertex $i_{1}$, as shown in Fig. 2. The product of any two $Z^{\otimes 3}$-type logical operators is a stabilizer because they must be representatives of a same logical operator. It then follows that $Z_{i_{j}}Z_{i_{j+1}}$ is a stabilizer for all $12\leq j\leq w+$ 1. Now, since $Z_{i_{2}}$ is not a stabilizer and must be an error operator that anti-commutes with some stabilizer generator, there is some stabilizer generator $g$ with $X$ or $Y$ acting on the qubit labeled by $i_{2}$. Because $g$ commutes with the stabilizer $Z_{i_{2}}Z_{i_{3}}$, it follows that $g$ has $X$ or $Y$ support on the qubit labeled by $i_{3}$ as well.
Figure 2: A subgraph of ${\mathfrak{G}}$ where vertex $i_{1}$ has an order greater than $w$. Here, the red edges correspond to the $Z^{\otimes 3}$-type logical operators. Since the product of two such logical operators is a stabilizer, it follows that $Z_{i_{j}}Z_{i_{j+1}}$ is a stabilizer for all $12\leq j\leq w+$ 1. We graphically denote these stabilizers by blue lines (note that these are not edges in ${\mathfrak{G}}$).
Similarly, we argue that $g$ must have $X$ or $Y$ support on the qubits labeled by $i_{4},i_{5},\ldots,i_{w+2}$. Then, $\operatorname{wt}(g)>w$, which is a contradiction. Hence, this case is not possible for $C$.
Case 2: Suppose that the order of each vertex is $\leq w$. Let $d_{1}$ be the number of logical $Z^{\otimes 3}$ operators with support on the first qubit, i.e., the number of edges in ${\mathfrak{G}}$. First, we will prove that if $V$ is a minimum vertex cover [^4] A vertex cover is a subset of the vertices such that at least one endpoint of each edge is in the subset. A minimum vertex cover is such a subset with the minimum number of vertices. of ${\mathfrak{G}}$, then the size of $V$ satisfies $|V|\geq\lfloor\frac{d_{1}}{2w}\rfloor$.
The matching of a graph is defined as a subset of its edges such that no two edges share a vertex. Let $M$ be a maximum matching of ${\mathfrak{G}}$. Due to the disjointness of the edges in $M$, each vertex in $V$ covers at most one edge of the maximum matching. Therefore, $|V|\geq|M|$. So, it is sufficient to prove that $|M|\geq\lfloor\frac{d_{1}}{2w}\rfloor$. Note that for $w=1$, the edges are disjoint; hence, $|M|=d_{1}\geq\lfloor\frac{d_{1}}{2w}\rfloor$. Now, suppose $w>1$. To generate an arbitrary matching, we consider the set of all edges. First, we pick a random edge from it and remove all the edges that share endpoints with it from the set. Since the highest order of the vertices is $\leq w$, the number of edges removed is at most $2(w-1)$. We choose another random edge from the set and continue this process until the set is a matching. In the end, we will obtain a matching of size $\geq\lfloor\frac{d_{1}}{2(w-1)+1}\rfloor\geq\lfloor\frac{d_{1}}{2w}\rfloor$. Hence, $|M|\geq\lfloor\frac{d_{1}}{2w}\rfloor$, which, in turn, implies that $|V|\geq\lfloor\frac{d_{1}}{2w}\rfloor$ for all $w$.
Now, the code must have a stabilizer generator $g$ that is $X$ or $Y$ on the first qubit because each $Z_{i}$ must anti-commute with some stabilizer generator. Suppose there is an edge between vertices labeled by $i$ and $j$, and so $Z_{1}Z_{i}Z_{j}$ is a logical operator. Since $g$ must commute with $Z_{1}Z_{i}Z_{j}$, it follows that $g$ has $X$ or $Y$ support on either the $i$-th or $j$-th qubit. Therefore, the vertices corresponding to the qubits that support $X$ or $Y$ of $g$ form a vertex cover of ${\mathfrak{G}}$. Hence, $\operatorname{wt}(g)\geq\lfloor\frac{d_{1}}{2w}\rfloor$. However, by the constant weight assumption, $w\geq\operatorname{wt}(g)$, which implies that $\lfloor\frac{d_{1}}{2w}\rfloor\leq w$. So, $d_{1}\leq 2w(w+1)$. Recall that $d_{1}$ is the number of logical $Z^{\otimes 3}$ operators with support on the first qubit (qubit with the highest degree). Hence, the total number of $Z^{\otimes 3}$-type logical operators satisfies $\ell\leq d_{1}n/3\leq 2w(w+1)n/3$. The first inequality here is from Lemma 2 that we will prove below separately. ∎
Here we prove a lemma where we establish a simple relation between the number of $Z^{\otimes 3}$-type logical operators and the degrees of qubits, which was used in the proof above and will also be used in later sections.
Lemma 2.
Suppose the number of $Z^{\otimes 3}$-type logical operators in an $n$-qubit stabilizer code is $\ell$. If $d_{i}$ is the degree of the $i$-th qubit of the code, then
$\max\{d_{i}\}_{i=1}^{n}\geq\frac{3\ell}{n}.$ (18)
Proof.
Recall that the degree of a qubit is defined to be the number of $Z^{\otimes 3}$-type logical operators that are supported on it. Without loss of generality, assume that the first qubit has the highest degree, i.e., $d_{1}=\max\{d_{i}\}_{i=1}^{n}$. The sum of the weights of all $Z^{\otimes 3}$-type logical operators is $3\ell$. This must be equal to the sum of the degrees of the qubits. So, $\sum_{i=1}^{n}d_{i}=3\ell$. But $\sum_{i=1}^{n}d_{i}\leq nd_{1}$. Hence, it follows that $nd_{1}\geq 3\ell$, and so $\max\{d_{i}\}_{i=1}^{n}\geq\frac{3\ell}{n}$. ∎
From Theorem 5 and Lemma 13, the best possible scaling of the estimation precision for the $ZZZ$ interaction using a stabilizer code with constant-weight stabilizer generators is the HS of $n^{-1}$. This family of stabilizer codes include quantum LDPC codes which are an important class of stabilizer codes as a candidate to achieve fault tolerance [38]. The thin surface code is one class of quantum LDPC codes, and by Theorem 1, it obtains the optimal scaling.
Theorem 5 also applies to cases where $w$ increases with respect to $n$. For example, one corollary of Theorem 5 is when $w=o(n)$, the optimal SHS $n^{-3}$ cannot be achieved. Note that it is also a corollary of Theorem 7 that we will present later in Sec. IV.3.
IV.2 Unachievability of the Optimal SHS
We now relax the constant-weight assumption on the stabilizer generators, in which case we want to establish a different no-go result that precludes a stabilizer code from achieving a precision scaling better than $1/n^{2}$.
A notable feature of the Shor code is its $Z^{\otimes 2}$-type stabilizer generators, as shown in Table 1. The following theorem demonstrates that the existence of such stabilizers is necessary for achieving the optimal SHS of $n^{-3}$.
Theorem 6.
Let $C$ be an $[[n,1]]$ stabilizer code. If $C$ does not have a $Z^{\otimes 2}$-type stabilizer, then the number of $Z^{\otimes 3}$-type logical operators is $\ell\leq{n(n-1)}/{6}=O(n^{2})$.
Proof.
Let $\mathcal{A}=\{A_{1},A_{2},\ldots A_{\ell}\}$ be the set of $Z^{\otimes 3}$-type logical operators. Let $d_{i}$ be the degree of the $i$-th qubit. Without loss of generality, we make the following three assumptions: (1) $d_{1}=\max\{d_{i}\}_{i=1}^{n}$, (2) $\mathcal{A}^{\prime}=\{A_{i}\}_{i=1}^{d_{1}}$ is the set of operators with support on the first qubit, and (3) The first $n^{\prime}$ qubits is the minimum set of qubits that contains the supports of operators in $\mathcal{A}^{\prime}$. All three assumptions can be satisfied by relabeling the qubits. The operators in $\mathcal{A}$ are different representatives of the same logical operator, and so the product of any two $A_{i},A_{j}\in\mathcal{A}^{\prime}$ is a stabilizer. If there is no $Z^{\otimes 2}$-type stabilizer, then each pair of operators in $\mathcal{A}^{\prime}$ only overlap on the first qubit. Hence, each operator in $\mathcal{A}^{\prime}$ is supported on two distinct qubits other than the first qubit. Therefore, $n^{\prime}=2d_{1}+1$. However, since $n^{\prime}\leq n$, we have $2d_{1}+1\leq n$. Combining this inequality with Eq. (18) in Lemma 2, we get
$\frac{3\ell}{n}\leq d_{1}\leq\frac{n-1}{2}.$ (19)
Therefore, $\ell\leq\frac{n(n-1)}{6}$, and so if there are no $Z^{\otimes 2}$-type stabilizers, then $\ell=O(n^{2})$. ∎
The scaling of QRM codes derived in Theorem 2 agrees with this result. One can show using the enumerator polynomial in Eq. (31) of Appx. B that QRM codes do not have any $Z^{\otimes 2}$-type stabilizer. Therefore, by Theorem 6 and Theorem 2, the QRM codes achieve the optimal precision scaling $1/n^{2}$ among all stabilizer codes without $Z^{\otimes 2}$-type stabilizers.
IV.3 Necessary Condition for Achieving the Optimal SHS
As we discussed, the existence of a $Z^{\otimes 2}$-type stabilizer is necessary for achieving the optimal SHS $\sim n^{-3}$. Note that the stabilizer group of Shor codes that achieve the optimal SHS contains a chain of $Z^{\otimes 2}$-type stabilizers. Motivated by that example, we provide a necessary condition for a stabilizer code that achieves this optimal SHS. Here, we prove that if the optimal SHS is achievable by a stabilizer code, then the stabilizer group contains such a chain of linear length.
Theorem 7.
Let $C_{n}$ be a family of $[[n,1,3]]$ stabilizer codes. Suppose that the number of $Z^{\otimes 3}$-type logical operators is $\ell=\Theta(n^{3})$. Then, for all $n$, there is a subset of qubits indexed by $\{i_{1},\ldots,i_{n_{r}}\}$ such that
  1. (1)
    $\{Z_{i_{1}}Z_{i_{2}},Z_{i_{2}}Z_{i_{3}},\ldots,Z_{i_{n_{r}-1}}Z_{i_{n_{r}}}\}$ is a subset of the stabilizers of $C_{n}$;
  2. (2)
    For any $1\leq k\leq r$, $Z_{i_{k}}$ is not a stabilizer of $C_{n}$;
  3. (3)
    $n_{r}=\Theta(n)$.
Proof.
Consider $C_{n}$. Let $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{\ell}\}$ be the set of $Z^{\otimes 3}$-type logical operators. Let $d_{i}$ denote the degree of the $i$-th qubit. Without loss of generality, we make the same three assumptions we made in Theorem 6: (1) $d_{1}=\max\{d_{i}\}_{i=1}^{n}$, (2) The set of operators with support on the first qubit is $\mathcal{A}^{\prime}=\{A_{i}\}_{i=1}^{d_{1}}$, and (3) The first $n^{\prime}$ qubits is the minimum set of qubits that contains the supports of operators in $\mathcal{A}^{\prime}$.
Define a graph ${\mathfrak{G}}$ with $(n^{\prime}-1)$ vertices $\{v_{2},v_{3},\ldots,v_{n^{\prime}}\}$ corresponding to the qubits indexed by $2,3,\ldots,n^{\prime}$, where an edge connects $v_{i}$ and $v_{j}$ if $Z_{1}Z_{i}Z_{j}\in\mathcal{A}^{\prime}$. In total, there are $d_{1}=|\mathcal{A}^{\prime}|$ edges in the graph. Since $\ell=\Theta(n^{3})$, Lemma 2 implies that $d_{1}=\Omega(n^{2})$. However, ${\mathfrak{G}}$ has $(n^{\prime}-1)$ vertices, and so it has at most $d_{1}=O(n^{\prime 2})=O(n^{2})$ edges. Thus, $d_{1}=\Theta(n^{2})$. Therefore, we can assume ${\mathfrak{G}}$ has $(n^{\prime}-1)\geq bn$ vertices with $d_{1}\geq an^{2}$ edges for some $a,b>0$ if we choose $n$ to be sufficiently large.
Note that the product of an even number of commuting logical operators ($Z^{\otimes 3}$-type operators here to be specific) in $\mathcal{A}$ is a stabilizer. Therefore, if there is an even path (a path consisting of an even number of edges) between $v_{i}$ and $v_{j}$ in ${\mathfrak{G}}$, then $Z_{i}Z_{j}$ is a stabilizer. We will show that we can always identify a subset of vertices of size $\Theta(n)$ such that there exists an even path between any pair of adjacent vertices.
By Lemma S1, ${\mathfrak{G}}$ has a connected component ${\mathfrak{G}}^{\prime}$ with $n_{1}^{\prime}=\Theta(n)$ vertices and $d_{1}^{\prime}=\Theta(n^{2})$ edges. Let the number of edges in ${\mathfrak{G}}^{\prime}$ be $d_{1}^{\prime}\geq(c/2)n_{1}^{\prime 2}$ for some constant $c\in(0,1)$. Suppose ${\mathfrak{G}}^{\prime}$ has $\geq c_{1}n_{1}^{\prime}$ vertices whose orders are $\leq c_{2}n_{1}^{\prime}$, where $c_{1},c_{2}\in(0,1)$. Then, the sum of the orders of the vertices is at most $c_{1}n_{1}^{\prime}\cdot c_{2}n_{1}^{\prime}+(1-c_{1})n_{1}^{\prime}\cdot n_{1}^{\prime}$. The sum of the orders of the vertices of a graph is twice the total number of edges. Therefore, $2\cdot(c/2)n_{1}^{\prime 2}\leq c_{1}c_{2}n_{1}^{\prime 2}+(1-c_{1})n_{1}^{\prime 2}$, and so
$c\leq 1-c_{1}+c_{1}c_{2}.$ (20)
Note that if we choose $c_{1}=1-\frac{c}{2}$ and $c_{2}=\frac{c}{3}$, we have $c\leq 1-c_{1}+c_{1}c_{2}=1-\left(1-\frac{c}{2}\right)+\left(1-\frac{c}{2}\right)\frac{c}{3}<\frac{c}{2}+\frac{c}{3}<c$, which is a contradiction. Therefore, ${\mathfrak{G}}^{\prime}$ has more than $(1-c_{1})n_{1}^{\prime}=\frac{c}{2}n_{1}^{\prime}$ vertices with orders greater than $\frac{c}{3}n_{1}^{\prime}$.
Reindex the vertices in descending order of their vertex orders, that is $v_{2}$ has the highest order, $v_{3}$ has the second-highest order and so on. Let $T$ be the set of the first $\lfloor\frac{c}{6}n_{1}^{\prime}\rfloor$ vertices. Each vertex in $T$ has an order $>\frac{c}{3}n_{1}^{\prime}$, which, in turn, is greater than $2|T|$. Therefore, for each vertex $v_{i}\in T$, we can assign a unique $\tilde{v}_{i}\not\in T$ that is connected to $v_{i}$ by a single edge such that $\tilde{v}_{i}\neq\tilde{v}_{j}$ for $i\neq j$. Now, starting from $i=2$, suppose that there is no even path between $v_{i}$ and $v_{i+1}$. Hence, all the paths from $v_{i}$ to $v_{i+1}$ consist of an odd number of edges. Then, by construction, there is an even path between $v_{i}$ and $\tilde{v}_{i+1}$, i.e., the path from $v_{i}$ to $v_{i+1}$ concatenated with the single edge from $v_{i+1}$ to $\tilde{v}_{i+1}$. We replace $v_{i+1}$ by $\tilde{v}_{i+1}$. Otherwise when there is already an even path between $v_{i}$ and $v_{i+1}$, we do no replacement. Similarly, we apply the same replacement rules for $i=3,4,\cdots,\lfloor\frac{c}{6}n_{1}^{\prime}\rfloor+1$. After all the appropriate replacements, all the adjacent vertices in $T$ are connected by even paths. Therefore, each pair of adjacent qubits support a $Z^{\otimes 2}$-type stabilizer. Furthermore, each such stabilizer is independent of the others. Hence, we conclude that, by reversing the reindexings, there exists a set of qubits labeled by $\{i_{1},i_{2},\ldots,i_{n_{r}}\}$ of size $n_{r}=\lfloor\frac{c}{6}n_{1}^{\prime}\rfloor=\Theta(n)$ such that $\{Z_{i_{1}}Z_{i_{2}},Z_{i_{2}}Z_{i_{3}},\ldots,Z_{i_{n_{r}-1}}Z_{i_{n_{r}}}\}$ is a subset of stabilizers.
From the discussion above, for any $1\leq k\leq r$, there exists some qubit $j_{k}$ such that $Z_{1}Z_{i_{k}}Z_{j_{k}}$ is a logical operator in $C_{n}$ because $Z_{i_{k}}$ must belong to some connected component of ${\mathfrak{G}}$. It implies $Z_{i_{k}}$ is not a stabilizer of $C_{n}$ because otherwise $Z_{1}Z_{j_{k}}=Z_{1}Z_{i_{k}}Z_{j_{k}}Z_{i_{k}}$ is a logical operator of $C_{n}$, contradicting with the fact that the code distance is $3$. ∎
Note that having $\{Z_{i_{1}}Z_{i_{2}},\ldots,Z_{i_{n_{r}-1}}Z_{i_{n_{r}}}\}$ as a part of the stabilizers of $C$ guarantees the codewords on qubits $\{i_{1},\ldots,i_{n_{r}}\}$ must be in the span of {|$\{\ket{0^{\otimes n_{r}}},\ket{1^{\otimes n_{r}}}\}$. Furthermore, as $\{Z_{i_{k}}\}_{1\leq k\leq r}$ are not stabilizers, for any choice of stabilizer generators there must be one of them that acts non-trivially as $X$ or $Y$ on all qubits in $\{i_{1},\ldots,i_{n_{r}}\}$. It implies the necessary existence of a weight-$\Theta(n)$ stabilizer generator when the optimal SHS is achieved. Theorem 7 shows the repetition code substructure described above is necessary in achieving the optimal SHS.
V Conclusion and outlook
In conclusion, we studied how different stabilizer codes perform when estimating the parameter $\omega$ encoded in the 3-local Hamiltonian in Eq. (2) under single-qubit noise. We also derived what properties stabilizer codes must satisfy in order to asymptotically achieve the SHS or the optimal SHS. Our work suggests more investigation towards the direction of exploiting the stabilizer codes for many-body quantum metrology.
The crux of the solution of our problem was the relation between the QFI and the number of logical $Z^{\otimes 3}$-type operators of the stabilizer code, which connects quantum metrology to the properties of stabilizer codes. Based on it, some of our results can be straightforwardly generalized to probing higher-order interactions under stronger quantum noise. For example, the $[[n,1,3]]$ thin surface code, QRM code and Shor code discussed in Sec. III all have natural generalizations to $[[n,1,d>3]]$ codes, that may be analogously applied in estimating higher-order interactions.
Many interesting questions remain open. For example, our results assumed the achievement of the HL, based on which we analyzed the achievability of the SHS. It would be interesting to consider whether better scalings with respect to the number of qubits will be available if we abandon the requirement of the HL. Another question is to ask how QEC protocols behave in many-body Hamiltonian estimation when taking fault-tolerance into consideration. Our current analysis was based on fast and precise QEC, and including measurement noise or circuit-level noise is left for future work.
Acknowledgments
S.Z. thanks Victor V. Albert and Michael Vasmer for helpful discussions. S.A. and S.Z. acknowledge funding provided by Perimeter Institute for Theoretical Physics, a research institute supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities.
Appendix A Optimal QFI coefficient
Here we compute the optimal QFI coefficient from Eq. (11) that provides the optimal coefficient of the HL among all possible QEC protocols. First, note that
$\quad{(\Delta G_{\mathrm{eff}})_{\mathrm{opt}}}/{2}=\min_{\text{2-local }S}\bigg{\|}\sum_{i<j<k}Z_{i}Z_{j}Z_{k}-S\bigg{\|}$ (21)
$=$ $\min_{\alpha_{ij}\!,\beta_{i}\!,\gamma\in{\mathbb{R}}}\bigg{\|}\!\sum_{i\!<\!j\!<\!k}\!Z_{i}Z_{j}Z_{k}-\!\sum_{i<j}\!\alpha_{ij}Z_{i}Z_{j}\!-\!\sum_{i}\!\beta_{i}Z_{i}\!-\!\gamma I\!\bigg{\|},$ (22)
because any off-diagonal term (in the computational basis) will only increase the operator norm. Next, let $A(\alpha,\beta,\gamma):=\sum_{i<j<k}Z_{i}Z_{j}Z_{k}-\sum_{i<j}\alpha_{ij}Z_{i}Z_{j}-\sum_{i}\beta_{i}Z_{i}-\gamma I$, where we use $\alpha$ (or $\beta$) to describe the column vector whose elements are $\alpha_{ij}$ (or $\beta_{i}$). We have $X^{\otimes n}A(\alpha,\beta,\gamma)X^{\otimes n}=-A(-\alpha,\beta,-\gamma)$. Thus, $\left\|2A(0,\beta,0)\right\|=\left\|A(\alpha,\beta,\gamma)+A(-\alpha,\beta,-\gamma)\right\|\leq\left\|A(\alpha,\beta,\gamma)\right\|+\left\|A(-\alpha,\beta,-\gamma)\right\|=2\left\|A(\alpha,\beta,\gamma)\right\|$, where we use the triangle inequality and the unitary invariance of operator norm. It implies that
${(\Delta G_{\mathrm{eff}})_{\mathrm{opt}}}/{2}=\min_{\beta\in{\mathbb{R}}^{n}}\left\|A(0,\beta,0)\right\|.$ (23)
Furthermore, let ${\mathcal{P}}$ represents the set of all permutations (of qubits) and $P$ its element, we have $\left\|A(0,{\mathrm{avg}}(\beta),0)\right\|\!\leq\!\frac{1}{\left|{\mathcal{P}}\right|}\!\sum_{P\in{\mathcal{P}}}\!\left\|A(0,P\beta,0)\right\|\!=\!\left\|A(0,\beta,0)\right\|$, where ${\mathrm{avg}}(\beta)=(\bar{\beta},\ldots,\bar{\beta})$, $\bar{\beta}=\frac{1}{n}\sum_{i}\beta_{i}$, $P\beta=(\beta_{P(1)},\beta_{P(2)},\ldots,\beta_{P(n)})$ and we use $\left\|A(0,P\beta,0)\right\|=\left\|A(0,\beta,0)\right\|$ for any permutation $P$. It implies
${(\Delta G_{\mathrm{eff}})_{\mathrm{opt}}}/{2}=\min_{\bar{\beta}\in{\mathbb{R}}}\left\|A(0,{\mathrm{avg}}(\beta),0)\right\|$ (24)
$=\min_{\bar{\beta}\in{\mathbb{R}}}\bigg{\|}\sum_{i<j<k}Z_{i}Z_{j}Z_{k}-\bar{\beta}\sum_{i}Z_{i}\bigg{\|}.$ (25)
For any computational state |$\ket{\phi(k)}$ that contains $k$ |$\ket{1}$s and $(n-k)$ |$\ket{0}$s, we have $A(0,{\mathrm{avg}}(\beta),0)\ket{\phi(k)}=\mu_{k}\ket{\phi(k)}$, where $\mu_{k}=\big{(}-\binom{k}{3}+\binom{k}{2}\binom{n-k}{1}-\binom{k}{1}\binom{n-k}{2}+\binom{n-k}{3}-\bar{\beta}(n-2k)\big{)}$. So,
$\mu_{k+1}-\mu_{k}=-4k^{2}+4(n-1)k+(2\bar{\beta}-(n-2)(n-1)).$ (26)
When $2\bar{\beta}\leq-(n-1)$, the value of $\mu_{k}$ increases as $k$ increases, and the maximum and minimum eigenvalues are taken at $k=0$ and $k=n$. When $2\bar{\beta}>-(n-1)$, there are four extreme points of $\mu_{k}$ taken at $k=0,n$ and locations near $\frac{(n-1)\pm\sqrt{n-1+2\bar{\beta}}}{2}$. Taking both situations and all extreme points into consideration, it is clear that the optimal $\bar{\beta}$ satisfies $2\bar{\beta}>-(n-1)$ and is taken at a point where the absolute values of the four extreme points are the same. As it is difficult to calculate the exact value of the optimal coefficient, we first conveniently assume $n-1$ is divisible by $4$, and take $\bar{\beta}=(n-1)^{2}/8-(n-1)/2$. In this case, the extreme points are taken at $k=0,\frac{n-1}{4},\frac{3(n-1)}{4},n$, and the corresponding eigenvalues of $A(0,{\mathrm{avg}}(\beta),0)$ are $\mu_{k}=\frac{(n+7)n(n-1)}{24},-\frac{(n+1)(n-1)(n-3)}{24},\frac{(n+1)(n-1)(n-3)}{24}$, $-\frac{(n+7)n(n-1)}{24}$. It implies
$\frac{(n+1)(n-1)(n-3)}{24}\leq\frac{(\Delta G_{\mathrm{eff}})_{\mathrm{opt}}}{2}\\ =\min_{\bar{\beta}\in{\mathbb{R}}}\bigg{\|}\sum_{i<j<k}Z_{i}Z_{j}Z_{k}-\bar{\beta}\sum_{i}Z_{i}\bigg{\|}\leq\frac{(n+7)n(n-1)}{24}.$ (27)
For general $n$, we have
${(\Delta G_{\mathrm{eff}})_{\mathrm{opt}}}=\frac{n^{3}}{12}+O(n^{2}).$ (28)
Appendix B Proofs of Properties of Shortened RM Codes
The classical RM codes have numerous interesting properties. In Sec. III.3, we outlined three key properties of the shortened RM code that are relevant to our discussion of QRM codes. Property (1), i.e., $\overline{RM}(1,m)\subsetneq\overline{RM}(m-2,m)$, is a simple observation since the rows of $\overline{{\mathsf{G}}}(1,m)$ are the same as the first $m$ rows of $\overline{{\mathsf{G}}}(m-2,m)$. Here, we prove Property (2) and (3).
Property 2.
The vector $\bm{1}_{2^{m}-1}$ and those corresponding to the rows of $\overline{{\mathsf{G}}}(m-2,m)$ form a basis for $\overline{RM}(1,m)^{\perp}$.
Proof.
We know that $RM(1,m)^{\perp}=RM(m-2,m)$ [55]. For the shortened code, we first delete the all-1 row and then delete the all-0 column. Hence, it immediately follows that $\overline{RM}(m-2,m)\subseteq\overline{RM}(1,m)^{\perp}$. Recall that $\dim\overline{RM}(1,m)=m=\binom{m}{m-1}$ and $\dim\overline{RM}(m-2,m)=\binom{m}{1}+\binom{m}{2}+\cdots+\binom{m}{m-2}$. Then, $\dim\overline{RM}(1,m)+\dim\overline{RM}(m-2,m)=2^{m}-2$, which is 1 less than the dimension of the column space. Hence, $\overline{RM}(1,m)^{\perp}$ has one extra basis vector that is not in the row space of $\overline{RM}(m-2,m)$. We claim that $\bm{1}_{2^{m}-1}$ is the additional basis vector for $\overline{RM}(1,m)^{\perp}$. Since the rows of $\overline{{\mathsf{G}}}(1,m)$ are of even weight, it is clear that $\bm{1}_{2^{m}-1}$ is orthogonal to each row vector, and so $\bm{1}_{2^{m}-1}\in\overline{RM}(1,m)^{\perp}$. It remains to show that $\bm{1}_{2^{m}-1}\not\in\overline{RM}(m-2,m)$.
For a contradiction, assume that $\bm{1}_{2^{m}-1}\in\overline{RM}(m-2,m)$. Then there are rows $\bm{v}_{i_{1}}^{T},\bm{v}_{i_{2}}^{T},\ldots\bm{v}_{i_{p}}^{T}$ of $\overline{{\mathsf{G}}}(m-2,m)$ such that $\bm{v}_{i_{1}}^{T}+\bm{v}_{i_{2}}^{T}+\cdots+\bm{v}_{i_{p}}^{T}=\bm{1}_{2^{m}-1}^{T}$. It follows that $(0|\bm{v}_{i_{1}}^{T})^{T}+(0|\bm{v}_{i_{2}}^{T})^{T}+\cdots+(0|\bm{v}_{i_{p}}^{T})^{T}=(0|\bm{1_{2^{m}-1}}^{T})^{T}\in RM(m-2,m)$. Adding $\bm{1}_{2^{m}}\in RM(m-2,m)$ to this equation, we obtain $\bm{1}_{2^{m}}+(0|\bm{v}_{i_{1}}^{T})^{T}+(0|\bm{v}_{i_{2}}^{T})^{T}+\cdots+(0|\bm{v}_{i_{p}}^{T})^{T}=(1|\bm{0}_{2^{m}-1}^{T})^{T}$, which is a weight-1 vector. However, $RM(m-2,m)$ has a minimum weight of $4$ (The minimum weight of $RM(r,m)$ is $2^{m-r}$ [55]). So, we have a contradiction; hence, property (2) is proven. ∎
Property 3.
The number of codewords in $\overline{RM}(1,m)^{\perp}$ with weight $(2^{m}-4)$ is $\frac{1}{6}(4^{m}-3\cdot 2^{m}+2)$.
Proof.
First, we consider the weight distribution, i.e., the number of codewords of each weight, of $RM(1,m)$. From that, we will obtain the weight distribution of $\overline{RM}(1,m)$, which, in turn, will give us the weight distribution of $\overline{RM}(1,m)^{\perp}$.
Let $A_{k}$ denote the number of codewords of weight $k$ in $RM(1,m)$. It can be shown that $A_{0}=1$, $A_{2^{m}}=1$, $A_{2^{m-1}}=2^{m+1}-2$, and $A_{k}=0$ for all $k\neq 0,1,2^{m-1}$ [55]. Note that adding the vector $\bm{1}_{2^{m}}^{T}$ to any weight-$(2^{m-1})$ vector of the form $(0|\bm{v})^{T}\in RM(1,m)$ results in another weight-$(2^{m-1})$ vector of the form $(1|\bm{v}^{\prime})^{T}$, where $\bm{v}$ and $\bm{v}^{\prime}$ are $(2^{m}-1)$-tuple row vectors. So, there is a one-to-one correspondence (related by the addition of $\bm{1}_{2^{m}}$) between the weight-$2^{m-1}$ vectors with the first entry 0 and the weight-$2^{m-1}$ vectors with the first entry 1. The shortening process gets rid of the all-1 row vector from the generator matrix. Therefore, the number of weight-$2^{m-1}$ vectors in $\overline{RM}(1,m)$ is half of that in $RM(1,m)$. Denoting $\overline{A}_{k}$ as the number of weight-$k$ codewords in $\overline{RM}(1,m)$, we get $\overline{A}_{0}=1$, $\overline{A}_{2^{m-1}}=2^{m}-1$ and $\overline{A}_{k}=0$ for all $k\neq 0,2^{m-1}$.
The enumerator polynomial $W(C;x,y)$ succinctly represents the weight distribution of a classical code $C$, with the coefficient of $x^{l}y^{n-l}$ representing the number of codewords of weight $l$. Based on the values of $\overline{A}_{k}$, the enumerator polynomial of $\overline{RM}(1,m)$ is
$W(\overline{RM}(1,m);x,y)=y^{2^{m}-1}+(2^{m}-1)x^{2^{m-1}}y^{2^{m-1}-1}.$ (29)
One can directly obtain the weight distribution of the dual code $C^{\perp}$ from the enumerator polynomial of $C$ using the Macwilliams identity
$W(C^{\perp};x,y)=\frac{1}{|C|}W(C;y-x,x+y),$ (30)
where $|C|$ is the number of codewords in $C$ [55]. Since $|\overline{RM}(1,m)|=2^{m}$, applying the Macwilliams identity to $W(\overline{RM}(1,m);x,y)$ yields
$W(\overline{RM}(1,m)^{\perp};x,y)$ $=\frac{1}{2^{m}}(x+y)^{2^{m}-1}$
$+\frac{(2^{m}-1)}{2^{m}}$ $(y-x)^{2^{m-1}}(x+y)^{2^{m-1}-1}.$ (31)
We want to find the coefficient of $x^{2^{m}-4}y^{3}$ on the right side of Eq. (31). Through a simple combinatorial calculation, the coefficient is found to be $\frac{1}{2^{m}}\binom{2^{m}-1}{3}+\frac{2^{m}-1}{2^{m}}\big{[}-\binom{2^{m-1}}{3}\binom{2^{m-1}-1}{0}+\binom{2^{m-1}}{2}\binom{2^{m-1}-1}{1}-\binom{2^{m-1}}{1}\binom{2^{m-1}-1}{2}+\binom{2^{m-1}}{0}\binom{2^{m-1}-1}{3}\big{]}$, which simplifies to $\frac{1}{6}(4^{m}-3\cdot 2^{m}+2)$. So, the number of weight-$(2^{m}-4)$ codewords in $\overline{RM}(1,m)^{\perp}$ is $\frac{1}{6}(4^{m}-3\cdot 2^{m}+2)$. ∎
Appendix C Graph Theory Lemma
In Sec. IV, we used several graph-theoretic tools to establish the no-go results. Specifically, in Sec. IV.3, we claimed that any graph with a quadratic number of edges has a connected component with a linear number of vertices and a quadratic number of edges. We now prove this claim in the following lemma.
Lemma S1.
Consider a family of graphs $\{{\mathfrak{G}}^{(n)}\}_{n}$ where $n$ belongs to an infinitely large set of positive integers. Assume ${\mathfrak{G}}^{(n)}$ has $\Theta(n)$ vertices and $\Theta(n^{2})$ edges. Then there is a family of connected components of ${\mathfrak{G}}^{(n)}$ with $\Theta(n)$ vertices and $\Theta(n^{2})$ edges.
Proof.
For a sufficiently large $n$, let ${\mathfrak{G}}={\mathfrak{G}}^{(n)}$ be a graph with $\geq an^{2}$ edges and $\geq bn$ vertices for some $a,b>0$. Let the connected components of ${\mathfrak{G}}$ be ${\mathfrak{G}}_{1}$, ${\mathfrak{G}}_{2}$, $\ldots,{\mathfrak{G}}_{k}$, and for all $i$, the number of vertices and edges in ${\mathfrak{G}}_{i}$ be $\bar{v}_{i}$ and $\bar{e}_{i}$, respectively. We index the connected components in descending order of the number of vertices. Let $\bar{e}=\max\{\bar{e}_{i}\}_{i=1}^{k}$. Suppose that the first $l$ connected components have at least $\bar{v}=\lceil\sqrt{2\bar{e}}+1\rceil$ vertices, i.e., enough vertices to make a complete graph with $\bar{e}$ edges.
Consider another graph ${\mathfrak{F}}$ with the same number of vertices as ${\mathfrak{G}}$ constructed in the following steps. Note that we will first partition vertices into different sets in steps (1) and (2) and then add edges within different sets in step (3).
  1. (1)
    Let $V_{1},V_{2},\ldots,V_{l}$ be subsets of vertices from ${\mathfrak{G}}_{1},\ldots,{\mathfrak{G}}_{l}$, where for each ${\mathfrak{G}}_{i}$ ($1\leq i\leq l$), we only select $\bar{v}$ vertices to place in $V_{i}$ and place the rest of the vertices in the set $V^{R}$.
  2. (2)
    Partition the set of vertices $V^{R}$ into subsets $V_{l+1},$ $V_{l+2},\ldots V_{k^{\prime}},V^{R^{\prime}}$ such that each $V_{i}$ ($1\leq i\leq k^{\prime}$) has $\bar{v}$ vertices and the remainder set $V^{R^{\prime}}$ has $\bar{w}<\bar{v}$ vertices.
  3. (3)
    Add edges between all vertices within each subsets of vertices $V_{1},V_{2},\ldots,V_{k^{\prime}},V^{R^{\prime}}$ to make them complete subgraphs ${\mathfrak{F}}_{1},{\mathfrak{F}}_{2},\ldots,{\mathfrak{F}}_{k^{\prime}},{\mathfrak{R}}$, which compose graph ${\mathfrak{F}}$.
Connected components ${\mathfrak{F}}_{1},{\mathfrak{F}}_{2},\ldots,{\mathfrak{F}}_{k^{\prime}}$ of ${\mathfrak{F}}$ have at least $\bar{e}$ edges each, and ${\mathfrak{R}}$ has $\bar{w}(\bar{w}-1)/2$ edges. So, the number of edges in ${\mathfrak{F}}$ is at least $\left(\frac{bn-\bar{w}}{\bar{v}}\right)\bar{e}+\frac{\bar{w}(\bar{w}-1)}{2}$. From the construction, ${\mathfrak{F}}$ has at least as many edges as ${\mathfrak{G}}$ does. Therefore,
$\left(\frac{bn-\bar{w}}{\sqrt{2\bar{e}}}\right)\bar{e}+\frac{\bar{w}(\bar{w}-1)}{2}\geq an^{2},$
$\Rightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ bn\sqrt{\bar{e}/2}+\sqrt{\bar{e}/2}+1\geq an^{2},$ (32)
where we set $\bar{w}=\sqrt{2\bar{e}}+2$ in the second step which increases the left-hand side so that the inequality still holds. Solving for $\bar{e}$, we get $\bar{e}=\Omega(n^{2})$. A graph with $\Omega(n^{2})$ edges must have $\Omega(n)$ vertices. However, the number of vertices in any connected component of ${\mathfrak{G}}$ is $O(n)$, which implies that $\bar{e}=O(n^{2})$. Therefore, $\bar{e}=\Theta(n^{2})$ and the number of vertices in a component with $\bar{e}$ edges is $\Theta(n)$. Hence, ${\mathfrak{G}}$ has a connected component with $\Theta(n)$ vertices and $\Theta(n^{2})$ edges. ∎